### v2

@@ 3.5 Task composability
In addition to facilitating accurate and reusable evaluations, inferring a task’s success from system state makes it easy to create composite tasks by combining together existing tasks. For example, the task “Create a calendar event with details and text the details to contact” could be created by combining together two existing tasks for creating a calendar event and for sending a text message, which is possible because each task initialization and success detection logic is hermetic. Composite tasks tend to be more challenging because of their complexity, although they provide partial rewards based on completion of sub tasks, to help facilitate hill climbing. The last two rows of Table 2 show the validation code for composite tasks.
@@ 3.6 Integrating MiniWob++
To demonstrate AndroidWorld’s extensibility, we implement MiniWoB++ in the AndroidWorld framework and term it MobileMiniWoB++. Each MobileMiniWoB++ task is instantiated using the standard AndroidWorld interface, inheriting from TaskEval base class, and contains methods like initialize_state and is_successful. Since MiniWoB++ leverages JavaScript for task configuration and success detection, we built a WebView app to communicate between Python and the app. For instance, the is_successful method of each task retrieves the reward value from the WebView app via an Android intent.
MobileMiniWoB++ introduces modifications in both observations and actions compared to the original benchmark. For example, HTML5 <input> elements are rendered with native Android UI widgets like the date-picker (see Figure  4), enhancing the realism of the tasks. MobileMiniWoB++ uses the same observation space as the Android tasks (accessibility tree and screenshot). Notably, it does not include the DOM as in the original implementation. The action space from AndroidWorld is retained. We manually review and test each task to ensure they are solvable. We excluded twelve of the original tasks that failed to render correctly on Android, presented compatibility issues with the touch interface, or required near real-time interaction, which poses challenges on emulators. Overall, AndroidWorld supports 92 MiniWoB++ tasks. See App. B.3 for more details.
@@ 4.1.1 M3A
After executing an action, M3A reflects on its effect by observing any state changes that may have occurred. During this stage, the agent is provided with available action types, general operating guidelines, the actual action taken, and its reasoning, as well as before-and-after UI states, represented by UI element representations and screenshots with SoM annotations. We request the agent to provide a concise summary of this step, including the intended action, success or failure, potential reasons for failure, and recommendations for subsequent actions. This summary will serve as the action history and be used for future action selection. See App. D for more details on the agent.
@@ 4.1.2 SeeAct baseline
To accommodate the Android environment, we adapt SeeAct in several ways. Firstly, we augment the action space from the original SeeAct implementation to support actions needed for mobile, including scroll, long press, navigate home and back, and open app actions. Secondly, in lieu of the DOM, which is not available for Android apps, we utilize the accessibility tree to construct candidate UI actions. Due to the lack of the DOM representation, we do not use the bespoke ranker model from the original implementation. However we observe that after applying a filtering heuristic to remove non-interactable elements, the majority of screens contains less than 50 candidate elements. See App. D.5 for more details on the implementation.
@@
Agent Input Base model SRAndroidWorld SRMMiniWoB++
@@ 4.2 Experiment results
Table 3 presents the success rates (SR) for the agents and human performance on both task suites. Although the agents have far from human performance, they demonstrate out-of-the-box capabilities in operating mobile UIs, exhibiting basic understanding and control capabilities of UIs. They can perform a variety of actions, including long-press, scrolling to search for information, and revising their plan if actions do not work out. The best performance is obtained for M3A when using GPT-4. On AndroidWorld the SoM-based variant is less performant, while on MobileMiniWoB++ it performs best. A similar result was obtained in recent work in the context of computer agents for desktop applications [65]. We posit SoM grounding is less beneficial for AndroidWorld as the accessibility tree for Android apps is generally better populated than for mobile websites and covers the required action space, reducing the value of SoM.
@@ 4.4 Agent robustness under random seeds
We evaluate agent robustness under two conditions: (1) identical tasks with the same parameters and (2) tasks with different parameter combinations, which change the initial state and task definition. Due to computational constraints, we perform this analysis on a representative subset of AndroidWorld tasks (listed in D.4). We use the strongest agent, M3A using the accessibility tree and GPT-4, for this analysis. Our results are shown in Figure 3.
Figure 3: Success rate variation across tasks due to the parametrization built into AndroidWorld. Using a fixed seed, the agent appears completely incapable of solving some tasks due to “bad luck” with the seed. In contrast, under different task parameterizations, we observe the agent is capable of solving the tasks fairly often. This contrast between the two groups is reflected by higher variance for the different seed group (orange) vs. the same seed group (blue). Significant differences, with p-value <  0.05, are indicated by “*”.
@@
Agent performance varies much more using different seeds, with statistically significant differences for the add expense and edit note tasks. The high intra-task variation indicates the model’s sensitivity to the seed and its parameters, stemming from a lack of robustness from the underlying model.
Research in RL environments ([20, 44, 9]) has noted similar sensitivity to seeds. Consistent with these studies, we observe agent performance is best represented by the mean across random seeds. Notably, the observation of non-zero rewards under some seeds points to potential enhancements through RL-like mechanisms in future work.
Overall, this experiment underscores the importance of extensively testing agents under varied task parameters to ensure they can handle real-world variability, a capability that AndroidWorld supports effectively.
@@ 5 Limitations
AndroidWorld currently supports tasks from mostly open-source Android apps with at least 1 million downloads and built-in Android system apps. While testing on more trending apps is desirable, we found open-source apps to be equally realistic and, in some cases, more challenging than apps with larger user bases. Trending apps tend to have UIs which are heavily optimized for smooth user experience, offering more UI functionality and shortcuts. Testing on less-optimized UIs makes the agent’s task harder. In an example failure we show in Figure 6(c), the agent needed to delete all notes in a list and failed by repeatedly searching for a “delete-all” button. Instead, an agent with stronger reasoning capabilities would have probably searched once for that functionality, realized it was not available, and deleted the notes one by one.
@@ C.4 Humans for task analysis
During development, we recruited six software engineers from Google to analyze task difficulty, duration, and category. Each human was assigned an equal portion of tasks and tasked with identifying bugs during this annotation phase. This process resulted in the discovery and resolution of over 30 bugs.

### v3

@@ 3.5 Task composability
In addition to facilitating accurate and reusable evaluations, inferring a task’s success from system state makes it easy to create composite tasks by combining together existing tasks. For example, the task “Create a calendar event with details and text the details to contact” was created by combining together two existing tasks for creating a calendar event and for sending a text message, which is possible because each task initialization and success detection logic is hermetic. Composite tasks tend to be more challenging because of their complexity, although they provide partial rewards based on completion of sub tasks, to help facilitate hill climbing. The last two rows of Table 2 show the validation code for composite tasks.
@@ 3.6 Integrating MiniWob++
We implement MiniWoB++ in the AndroidWorld framework and term it MobileMiniWoB++. Each MobileMiniWoB++ task is instantiated using the standard AndroidWorld interface, inheriting from TaskEval base class, and contains methods like initialize_state and is_successful. Since MiniWoB++ leverages JavaScript for task configuration and success detection, we built a WebView app to communicate between Python and the app. For instance, the is_successful method of each task retrieves the reward value from the WebView app via an Android intent.
MobileMiniWoB++ introduces modifications in both observations and actions compared to the original benchmark. For example, HTML5 <input> elements are rendered with native Android UI widgets like the date-picker (see Figure  4), enhancing the realism of the tasks. MobileMiniWoB++ uses the same observation space as the Android tasks (accessibility tree and screenshot). Notably, it does not include the DOM as in the original implementation. The action space from AndroidWorld is retained. We manually review and test each task to ensure they are solvable. We excluded twelve of the original tasks that failed to render correctly on Android, presented compatibility issues with the touch interface, or required near real-time interaction, which poses challenges on emulators. Overall, AndroidWorld supports 92 MiniWoB++ tasks. See Appendix B.3 for more details.
@@ 4.1.1 M3A
After executing an action, M3A reflects on its effect by observing any state changes that may have occurred. During this stage, the agent is provided with available action types, general operating guidelines, the actual action taken, and its reasoning, as well as before-and-after UI states, represented by UI element representations and screenshots with SoM annotations. We request the agent to provide a concise summary of this step, including the intended action, success or failure, potential reasons for failure, and recommendations for subsequent actions. This summary will serve as the action history and be used for future action selection. See Appendix D for more details on the agent.
@@ 4.1.2 SeeAct baseline
To accommodate the Android environment, we adapt SeeAct in several ways. Firstly, we augment the action space from the original SeeAct implementation to support actions needed for mobile, including scroll, long press, navigate home and back, and open app actions. Secondly, in lieu of the DOM, which is not available for Android apps, we utilize the accessibility tree to construct candidate UI actions. Due to the lack of the DOM representation, we do not use the bespoke ranker model from the original implementation. However, we observe that after applying a filtering heuristic to remove non-interactable elements, the majority of screens contains less than 50 candidate elements. See Appendix D.5 for more details on the implementation.
@@
Agent Input Base model SRAndroidWorld SRMobileMiniWoB++
@@ 4.2 Experimental results
Table 3 presents the success rates (SR) for the agents and human performance on both task suites. Although the agents have far from human performance, they demonstrate out-of-the-box capabilities in operating mobile UIs, exhibiting basic understanding and control capabilities of UIs. They can perform a variety of actions, including long-press, scrolling to search for information, and revising their plan if actions do not work out. The best performance is obtained for M3A when using GPT-4. On AndroidWorld the SoM-based variant is less performant, while on MobileMiniWoB++ it performs best. A similar result was obtained in recent work in the context of computer agents for desktop applications [66]. We posit SoM plays a more critical role in MobileMiniWoB++ tasks due to the often incomplete accessibility tree, compared to that of native Android apps.
@@ 4.4 Agent robustness under random seeds
We evaluate agent robustness under two conditions: (1) identical tasks with the same parameters and (2) tasks with different parameter combinations, which change the initial state and task definition. Due to computational constraints, we perform this analysis on a representative subset of AndroidWorld tasks (listed in Appendix D.4). We use the strongest agent, M3A using the accessibility tree and GPT-4, for this analysis. Our results are shown in Figure 3.
Figure 3: Success rate variation across tasks due to the parametrization built into AndroidWorld. Using a fixed seed, the agent appears completely incapable of solving some tasks due to “bad luck” with the seed. In contrast, under different task parameterizations, we observe the agent is capable of solving the tasks fairly often. Wilson binomial proportion confidence intervals (95%) are shown for the different seed group (orange) and the same seed group (blue). The different seed group has higher variance than the same seed group. Significant differences, with p-value <  0.05, are indicated by “*”.
@@
Agent performance varies much more using different seeds, with statistically significant differences for the add expense and edit note tasks. The high intra-task variation indicates the model’s sensitivity to the seed and its task parameters, stemming from a lack of robustness from the underlying model.
Research in RL environments [21, 44, 10] has noted similar sensitivity to seeds. Consistent with these studies, we observe agent performance is best represented by the mean across random seeds. Notably, the observation of non-zero rewards under some seeds points to potential enhancements through RL-like mechanisms in future work.
This experiment underscores the importance of extensively testing agents under varied task parameters to ensure they can handle real-world variability, a capability that AndroidWorld supports effectively.
@@ 5 Limitations
AndroidWorld currently supports tasks from open-source Android apps with at least 1 million downloads and from built-in Android system apps. While testing on more trending apps is desirable, we found open-source apps to be equally realistic and, in some cases, more challenging than apps with larger user bases. Trending apps tend to have UIs which are heavily optimized for smooth user experience, offering more UI functionality and shortcuts. Testing on less-optimized UIs makes the agent’s task harder. In an example failure we show in Figure 6(c), the agent needed to delete all notes in a list and failed by repeatedly searching for a “delete-all” button. Instead, an agent with stronger reasoning capabilities would have probably searched once for that functionality, realized it was not available, and deleted the notes one by one.
@@ C.4 Humans for task analysis
During development, we recruited six volunteers with proficient programming skills to analyze task difficulty, duration, and category. Each human was assigned an equal portion of tasks and tasked with identifying bugs during this annotation phase. This process resulted in the discovery and resolution of over 30 bugs.